# Adaptive Reduced Multilevel Splitting

### v1

@@ Abstract
This paper considers the classical problem of sampling with Monte Carlo methods a target rare event distribution of the form $\eta^{\ast}_{l_{\mathrm{max}}}\propto\,{\rm l}{\rm l}_{S^{\ast}>l_{\mathrm{max}}}d\pi$, with $\pi$ a reference probability distribution, $S^{\ast}$ a real-valued function that is very expensive to compute and $l_{\mathrm{max}}\in\mathbb{R}$ some level of interest. We assume we can iteratively build a sequence of reduced models $\{S^{(k)}\}_{k}$ of the scores, which are increasingly refined approximations of $S^{\ast}$ as new score values are computed; these reduced models being moreover certified with error bounds. This work proposes a fully adaptive algorithm to iteratively build a sequence of proposal distributions with increasing target levels $l^{(k)}$ of the form $\propto\,{\rm l}{\rm l}_{S^{(k)}>l^{(k)}}d\pi$ and draw from them, the objective being importance sampling of the target rare events, as well as proposing relevant updates to the reduced score. An essential contribution consists in adapting the target level to the reduced score: the latter is defined as the first non-achievable level $l^{(k)}$ corresponding to a cost, in a pessimistic scenario, for importance sampling matching an acceptable budget; it is calculated solely from the reduced score function and its error bound. From a practical point of view, sampling the proposal sequence is performed by extending the framework of the popular adaptive multilevel splitting algorithm to the use of reduced score approximations. Numerical experiments evaluate the proposed importance sampling algorithm in terms of computational complexity versus squared error. In particular, we investigate the performance of the algorithm when simulating rare events related to the solution of a parametric PDE, which is approximated by a reduced basis.
@@
Rare event simulation, reduced modeling, importance sampling, adaptive multilevel splitting.
@@ 1 Introduction
1.1 Context
Let $\left(\eta^{\ast}_{l}\right)_{l\in\mathbb{R}}$ denote a family of target probability distributions defined on a state space $\mathcal{X}$ by the conditional probability distributions
$\eta^{\ast}_{l}\displaystyle\mathop{=}^{{\rm def}}\frac{1}{p_{l}^{\ast}}\,{\rm l}{\rm l}_{S^{\ast}>l}\mathrm{d}\pi\qquad l\in\mathbb{R}.$ (1.1)
In the above, $\pi$ denotes a given, easily simulable, reference probability distribution. We have also introduced a real-valued score function
$S^{\ast}:\mathcal{X}\to\mathbb{R},$
which is assumed to be of the form
@@ 1.1 Context
where $\mathrm{score}$ is an easily computable function of physical interest, and $\Psi^{\ast}(x)$ is the outcome of a presumably exact numerical computation of a complex physical system parametrized by $x\in\mathcal{X}$. For instance, the random variable $x$ can describe unknown physical parameters, uncertainty on initial/boundary conditions, or additional noise modeling either a chaotic behavior or the interaction with an environment.
@@
The quantity $p^{\ast}_{l}\displaystyle\mathop{=}^{{\rm def}}\int_{\mathcal{X}}\,{\rm l}{\rm l}_{S^{\ast}>l}\mathrm{d}\pi=\pi\left(\left\{S^{\ast}>l\right\}\right)$ is the normalization of interest, here the probability of the considered rare events $\left\{S^{\ast}>l\right\}$, for $l\in\mathbb{R}$.
@@
We consider the classical problem of designing an algorithmic Monte Carlo procedure that samples according to $\eta^{\ast}_{\mathrm{max}}$ – for one specific large $l_{\mathrm{max}}\in\mathbb{R}$ – and estimate the associated rare event probability $p^{\ast}_{l_{\mathrm{max}}}$ (the normalization).
In this work, we will in addition be interested in scenarios in which the practitioner has to face two types of difficulty.
@@
The first difficulty is classical –especially in high dimension–, and arises when $l$ increases, yielding a rare event (or low temperature problem). A rare event problem is defined by a very small yet strictly positive probability $p^{\ast}_{l}$; the analogue in statistical mechanics being the low temperature regime111for instance when considering the Gibbs (canonical) distribution proportional to $\mathrm{e}^{lS^{\ast}}\mathrm{d}\pi$.. Both settings confronts the practitioner with similar computational challenges: the density defining $\eta_{l}^{\ast}$ concentrates the majority of the probability mass in specific areas of $\mathcal{X}$, which are roughly described by the maxima of $S^{\ast}$. A direct sampling with a Monte Carlo simulation of $\pi$ requires a sample size of order at least $1/p^{\ast}_{l}$, which is often infeasible. In that context, a very popular general strategy to simulate a sample according to $\eta^{\ast}_{l}$ and to estimate the associated normalizing constant is to resort to a Sequential Monte Carlo (SMC) strategy (see [10]) that starts with a Monte Carlo sample of particles distributed according to $\pi$ and then sequentially samples $\eta^{\ast}_{l}$ for increasing values of $l$ by resorting to a combination of Importance Sampling (IS), re-sampling (selection) of weighted samples and mutations of particles based on suitable Markov Chain Monte Carlo transitions. More specifically, we are interested in an adaptive variant designed for the rare event context, namely the Adaptive Multilevel Splitting (AMS) algorithm (see [5]). Some reader may find convenient to note that AMS is a close relative to subset simulation [1].
@@
The second difficulty we will consider here appears when the numerical evaluation of Ψsuperscriptnormal-$\Psi^{\ast}$ is extremely expensive; so expensive that the number of evaluations required in, say, a usual SMC approach becomes prohibitive. In order to circumvent this issue, we assume that the physical model $\Psi^{\ast}$ can be approximated by a reduced model denoted $\Psi$; the cost of evaluating $\Psi$ being small as compared to one evaluation of the true score with $\Psi^{\ast}$. The reduced models $\Psi$ we consider is assumed to come with two key features. To explain these, let us denote from now on the approximate score function by
@@
The first feature is an a posteriori pointwise error quantification associated with $S$. This error quantification comes in the form of an error bound denoted
$E(x)=\mathrm{error}\left(\Psi(x)\right)\in\mathbb{R},$
and satisfying the pointwise estimate:
@@
which takes as an input any sequence of states in $\mathcal{X}$ – called here a sample of snapshots – together with the solution of the associated true model $\Psi^{\ast}$ evaluated for each of those snapshots. The output is the reduced model $\Psi$. In order to improve the readability, and throughout the paper, we may use the superscript $(k)$ instead of $(X_{1},\ldots,X_{k})$, or even drop it, as follows:
$(S^{(X_{1},\ldots,X_{k})},E^{(X_{1},\ldots,X_{k})})=(S^{(k)},E^{(k)})=(S,E),$
and we will proceed equivalently for any quantity that depends on the model reduction procedure through $(X_{1},\ldots,X_{k})$, the sample of snapshots.
A reduced model can be constructed by various, potentially very different, means. In this work, we will consider the reduced modeling procedure and its error quantification satisfying (1.2) as a black box. However, the numerical tests will be performed with two specific reduced modeling approach. The first test, based on splines, is artificially challenging (for the simulation method) and is not meant to be used in practice. The second test uses a reduced basis model reduction. Both rigorously satisfy  (1.2).
@@ 1.2 Previous works
Reduced modeling
The idea of using a simpler surrogate model in Monte-Carlo simulation for target distributions is not new and there is a large amount of literature on the subject. Most references are concerned with either kriging on the one hand, and reduced basis methods on the other hand.
A popular Bayesian approach is given by kriging, in which the surrogate score function $S:\mathcal{X}\to\mathbb{R}$ is a random field, distributed according to the posterior of a prior Gaussian field conditioned by fitting the exact values of the evaluated snapshots. The estimate (1.2) is however not satisfied almost surely, and one has to resort to averaged error estimates.
Among deterministic approaches, we will especially be interested in reduced basis methods, which are known to be very efficient for the numerical approximation of problems involving the repeated solution of parametric Partial Differential Equations (PDEs). The strategy adopted consists in projecting the associated high dimensional numerical problem on a subspace where the solutions corresponding to suitably chosen parameters are efficiently approximated and characterized with precise error bounds of the form (1.2). An overview on the theoretical rationale behind reduced basis and the practical application of this method is provided in the reference book [21].
The list of references using kriging is the following: [11, 19, 22, 25, 2, 17, 3], while those using reduced basis is [8, 16, 13, 15]. Generic or other meta-models are considered in [4, 12, 24, 20]; the special problem of well-tuning a combination of possibly generic multiresolution or multifidelity surrogates is treated in [12, 24], as well as in the survey [20].
Snapshot sampling
@@
Arguably, the most sensitive feature of such reduced model approach to sampling is the specific collection of snapshots (with associated evaluation of the true model) $(X_{k},\Psi^{\ast}(X_{k}))$, $k\geq 1$ with which the reduced model is updated after each iteration $k$. A natural adaptive approach in this context consists in sampling sequentially the collection of snapshots, the reduced model being updated after each iteration $k$. The updated reduced model (and its associated error estimates) can then be used to pick the next snapshot in presumably interesting areas.
@@
The sampling or selection of snapshot $X_{k}$ then follows from two ingredients: i) a probability distribution used to generate candidates, and ii) a weighting or fitness function (sometimes called learning function) used to pick the presumably best candidate. The idea of ingredient ii) can be found in the seminal work of [11] where selecting snapshots is called active learning. In the latter approach, snapshots are selected by optimizing a criterion (the learning function) that will most likely reduce the uncertainty of the kriging surrogate model. An exhaustive set of references and a very quantitative review of methods related to active learning and kriging is provided in [19]. Some other references, as [16], pick the snapshots $(X_{k})_{k\geq 1}$ uniformly among candidates using a decreasing sequence of subsets that contain the target rare event, constructed using the iteratively updated error estimates.
@@
As we will argue in the present work, ingredient i) is also of high importance. Indeed, a key general phenomenon in a rare-event and high-dimensional setting is the following: the true target distribution $\eta^{\ast}_{l}$ is extremely concentrated around very specific but unknown points in $\mathcal{X}$, so that, unless the pool of candidates among which snapshots $(X_{1},\ldots,X_{k})$ are picked is partly also located in the same regions, the information given by the reduced model $\Psi^{(X_{1},\ldots,X_{k})}$ may not be accurately updated. As a consequence, unless one is ready to sample an unreasonable amount of candidates used to optimize the learning function, it is critical to select such candidates in areas that carefully “approach” the true rare event. Note also that sampling the snapshot with a well-identified distribution (instead of an optimization procedure) enables to perform importance sampling with the exact model $S^{\ast}$; an approach we will use in the present work.
Most references in the literature (at least those cited here) construct pools of candidates for the choice of snapshot $X_{k}$ at iteration $k$ by relying on either the reference distribution $\pi$ (in an offline or preliminary phase), or (in some way or another) on the history of states resulting from sampling at iteration $k$ surrogates of the target conditional distributions (1.1) built using the reduced score, defined by
$\mu_{l}\displaystyle\mathop{=}^{{\rm def}}\frac{1}{p_{l}}\,{\rm l}{\rm l}_{S>l}\mathrm{d}\pi\qquad l\in\mathbb{R}.$
Again, a potential issue is that for large $l$, the reduced rare event is also rare, hence very concentrated, and thus usually off, being too far away from the true target rare event. Selecting carefully an appropriate $l$ by leveraging the existence of error estimates is a key novelty introduced in the present work. This delicate issue is addressed in very few papers in the literature. In a reverse approach, let us mention the worst-case error criterion proposed in [15], which is used to select the reduced score $S$ in a hierarchy of precomputed surrogates, for each level $l$ of a cross-entropy optimization process.
Sampling the reduced rare event
The general framework discussed above has been used extensively for rare event simulation in conjunction with several Monte Carlo strategies used to sample the reduced conditional distributions $\mu_{l}$. Sampling strategies can be grouped in two broad categories: on the one hand importance sampling (IS), usually through a parametric approach (a popular advanced method being e.g. adaptive IS with cross entropy minimization), and on the other hand split/ subset / adpative multilevel splitting (AMS) / sequential Monte Carlo (SMC) simulation methods which are non-parametric (albeit targeting an approximate, surrogate distribution).
Let us present a brief review of recent contributions that exploit this general framework of adaptive surrogate modeling for rare event simulation. We will consider IS approaches on the one hand and AMS simulation on the other.
Many adaptive approaches to importance sampling have been proposed in recent years. In particular, a popular criterion for finding a good proposal in a parametric family is obtained by minimizing the relative entropy between the target and the proposal, resulting in the cross-entropy method ([9]). This method has been proposed for surrogate targets based on kriging in [2], or reduced basis approximations in [13, 15]. Other recent contributions also builds adaptive proposals for IS using kriging models [22, 25]. In order to increase the effectiveness of the cross-entropy method, let us mention that it has been suggested to approximate the parameter variable $X$ by constraining its support in well-chosen low dimensional subspaces of $\mathcal{X}$ [23]. Stratified sampling (that can be interpreted as a form of IS) with a surrogate model is developed in [4].
Similarly to adaptive IS, sampling with splitting methods (subset simulation, AMS) have been used in conjunction with kriging-based surrogate models (constructed adaptively) in [17, 3]. We refer again to the review on methods related to active learning and kriging provided in [19]. In [12] subset simulation is performed using nested multiresolution sets (provided by an accuracy hierarchy of the surrogate model), yielding an estimate with an associated a posteriori error. Instead of targeting distributions defined by conditioning on surrogate rare events, one can also use sequential Monte Carlo with soft levels and tempering. This is the approach chosen in [24] with a score function also relying on a multiresolution approximation of a PDE solution.
Adapting splitting methods to reduced scores generally requires the definition of a bridging procedure. The goal of this bridge is to sample a new proposal $\mu^{\prime}_{l^{\prime}}$ defined with some updated reduced score ${S}^{\prime}$ and level $\l^{\prime}$, starting from the samples of the current proposal $\mu_{l}$. In general, this sampling is not trivial. For example, the nesting underlying an AMS algorithm is broken, since in general updating the reduced score and the level will imply $\left\{{S}^{\prime}>l^{\prime}\right\}\not\subset\left\{S>l\right\}$. There is very little work in the literature on this bridging problem. However, we would like to mention the work [24], which proposes a bridging procedure in the context of subset simulation with a level $l^{\prime}=l$ maintained along the bridge.
1.3 Sketch of the proposed method
In this work, we will consider the surrogates of the target conditional distributions (1.1) built using the reduced score
@@ 1.3 Sketch of the proposed method
as proposal distributions, where the normalization $Z_{l}\displaystyle\mathop{=}^{{\rm def}}\int_{\mathcal{X}}\,{\rm l}{\rm l}_{S>l}\mathrm{d}\pi,$ represents an approximation of the probability of the rare event. The main iteration index of the main algorithm is $k$, the total number of sampled snapshots already evaluated with the true model. The main algorithm is basically the following.

### v3

@@ Abstract
This paper considers the classical problem of sampling with Monte Carlo methods a target rare event distribution defined by a score function that is very expensive to compute. We assume we can build using evaluations of the true score, an approximate surrogate score certified with error bounds. This work proposes a fully adaptive algorithm to sequentially sample surrogate rare event distributions with increasing target levels. An essential contribution consists in sampling at each iteration the surrogate rare event at a critical level corresponding to a specific cost. This cost is related to importance sampling for a target for a given budget. The critical level is calculated solely from the reduced score and its error bound. From a practical point of view, sampling the proposal sequence is performed by extending the framework of the popular adaptive multilevel splitting algorithm to the use of score approximations. Numerical experiments evaluate the proposed importance sampling algorithm in terms of computational complexity versus squared error. In particular, we investigate the performance of the algorithm when simulating rare events related to the solution of a parametric PDE, which is approximated by a reduced basis.
@@
Rare event simulation, reduced modeling, importance sampling, adaptive multilevel splitting, relative entropy.
@@ 1 Introduction
1.1 Context and objective
Let $\pi(\mathrm{d}x)$ denotes a given, easily simulable, reference probability distribution on a state space $\mathcal{X}$. A random variable with distribution $\pi$ may describe for instance unknown physical parameters, uncertainty on initial/boundary conditions, or additional noise modeling either a chaotic behavior or the interaction with an environment. We also consider a real-valued score function $S^{\ast}:\mathcal{X}\to\mathbb{R}$ which is assumed to be of the form
@@ 1.1 Context and objective
where $\mathrm{score}$ is an easily computable function of physical interest, and $\Psi^{\ast}(x)$ is the outcome of a presumably exact numerical computation of a complex physical system parametrized by $x\in\mathcal{X}$. With the previous ingredients, one can define the following family of conditional probability distributions on $\mathcal{X}$:
$\eta^{\ast}_{l}(\mathrm{d}x)\displaystyle\mathop{=}^{{\rm def}}\frac{1}{p_{l}^{\ast}}\,{\rm l}{\rm l}_{S^{\ast}(x)>l}\pi(\mathrm{d}x)\qquad l\in\mathbb{R}$ (1.1)
@@
defined by conditioning the distribution $\pi$ on the event $\left\{S^{\ast}>l\right\}$ which may be interpreted as a certain failure region/event.The quantity $p^{\ast}_{l}\displaystyle\mathop{=}^{{\rm def}}\int_{\mathcal{X}}\,{\rm l}{\rm l}_{S^{\ast}>l}\mathrm{d}\pi=\pi\left(\left\{S^{\ast}>l\right\}\right)$ is the associated normalization, here the probability of the considered (possibly rare) event $\left\{S^{\ast}>l\right\}$, for $l\in\mathbb{R}$. $\eta_{l}^{\ast}$ is the probability distribution of the system conditioned to the failure event.
@@
We consider in this paper the classical problem of designing an algorithmic Monte Carlo procedure that samples according to $\eta^{\ast}_{l_{\mathrm{max}}}$ – for one specific large $l_{\mathrm{max}}\in\mathbb{R}$ – and estimate the associated rare event probability $p^{\ast}_{l_{\mathrm{max}}}$. Note that depending on the problem at hand, the final event of interest might be either $\left\{S^{\ast}>l_{\mathrm{max}}\right\}$ or $\left\{S^{\ast}\geq l_{\mathrm{max}}\right\}$, the intermediate sets still being defined as $\left\{S^{\ast}>l\right\}$ for $l<l_{\mathrm{max}}$.
@@
In this work, we will focus on scenarios with two types of difficulty. A first difficulty arises when ${l_{\mathrm{max}}}$ is large, generating a very small (yet strictly positive) probability $p^{\ast}_{l_{\mathrm{max}}}>0$. This rare event context confronts the practitioner with a first computational challenge: the majority of the probability mass of the conditional distribution $\eta_{l_{\mathrm{max}}}^{\ast}$ is concentrated in specific areas of $\mathcal{X}$, which are roughly described by the maxima of $S^{\ast}$ which are usually unknown. A direct sampling with a Monte Carlo simulation of $\pi$ requires a sample size of order at least $1/p^{\ast}_{l_{\mathrm{max}}}$, which is often infeasible. In that context, a popular efficient strategy to simulate a sample according to $\eta^{\ast}_{l_{\mathrm{max}}}$ and to estimate the associated normalizing constant is to resort to ’Sequential Monte Carlo’ (SMC) methods. Such methods start with a Monte Carlo sample of clones (also called below ’particles’) of the system distributed according to $\pi$ and then sequentially samples $\eta^{\ast}_{l}$ for increasing values of $l$. To do so, those methods resort to a combination of Importance Sampling (IS), re-sampling (selection) of weighted samples, and mutations of particles based on suitable Markov Chain Monte Carlo transitions. In this work we will focus on a specific instance called Adaptive Multilevel Splitting (AMS, see [5, 6]) which moreover chooses adaptively the sequence of levels $l$ where re-sampling is performed.
@@
A second difficulty appears when the numerical evaluation of $\Psi^{\ast}$ is extremely expensive, so that the number of evaluations required in, say, a usual SMC or AMS approach becomes prohibitive. In order to circumvent this issue, we assume that the physical model $\Psi^{\ast}$ can be approximated by a reduced model denoted $\Psi$, the cost of evaluating $\Psi$ being small as compared to one evaluation of the true score with $\Psi^{\ast}$. The reduced model $\Psi$ we consider is assumed to come with two key features. To explain these, let us denote from now on the approximate score function by
@@
The first feature is an a posteriori pointwise error quantification associated with $S$. This error quantification comes in the form of an explicit error bound denoted $E(x)=\mathrm{error}\left(\Psi(x)\right)\in\mathbb{R},$ and satisfying the pointwise estimate:
@@
which takes as an input any sequence of states in $\mathcal{X}$ – called here a sample of snapshots – together with the solution of the associated true model $\Psi^{\ast}$ evaluated for each of those snapshots. The output is the reduced model $\Psi$. To improve the readability of the document, we will use a superscript $(k)$ to specify the dependency on $(X_{1},\ldots,X_{k})$, or even omit this dependency if the context is unambiguous,
$(S^{(X_{1},\ldots,X_{k})},E^{(X_{1},\ldots,X_{k})})=(S^{(k)},E^{(k)})=(S,E).$
@@ 1.2 Previous works
Reduced models for rare events
The idea of using a simpler surrogate model in Monte-Carlo simulation for target distributions is not new and there is a large amount of literature on the subject. Most references are concerned with either kriging on the one hand, and reduced basis methods on the other hand. For kriging approaches [12, 20, 23, 27, 2, 18, 3], the surrogate score function $S:\mathcal{X}\to\mathbb{R}$ is a random field, distributed according to the posterior of a prior Gaussian field conditioned by fitting the exact values of the evaluated snapshots. The estimate (1.2) is however not satisfied almost surely, and one has to resort to averaged error estimates. For reduced basis approaches [9, 17, 14, 16], the surrogate score function is deterministic, and constructed by projection of the function $\Psi^{\ast}$ on a subspace where it is efficiently approximated and characterized with precise error bounds of the form (1.2). This model reduction framework is known to be very efficient for the numerical approximation of problems involving the repeated solution of parametric Partial Differential Equations (PDEs) [22]. We note that besides these two main approaches, other families of reduced models [4, 13, 26, 21] or combination of them [13, 26, 21] have been proposed for rare event simulation.
@@
Arguably, the most sensitive feature of any of these reduced model approaches is the specific collection of snapshots $(X_{k},\Psi^{\ast}(X_{k}))$, $k\geq 1$ with which the reduced model is updated along the iterations. A natural approach consists in sampling sequentially and adaptively the collection of snapshots, the reduced model being updated after each new evaluation of the true model at iteration $k$. The updated reduced model (and the associated error estimates) can then be used to pick the next snapshot in presumably interesting areas. Sampling the snapshot $X_{k}$ rely on two ingredients: i) a probability distribution used to generate candidates, and ii) a weighting function, also called learning function, used to pick the presumably best candidate.
@@
The idea of ingredient ii) can be found in the seminal work of [12] where selecting snapshots is called active learning. In the latter approach, snapshots are selected by optimizing a criterion (the learning function) that will most likely reduce the uncertainty of the surrogate model. An exhaustive set of references and a very quantitative review of methods related to active learning and kriging is provided in [20]. Some other references, as [17], pick the snapshots $(X_{k})_{k\geq 1}$ uniformly among candidates using a decreasing sequence of subsets that contain the target rare event, constructed using the iteratively updated error estimates.
@@
However, as we will argue in the present work, ingredient i) is also of high importance. Indeed, a key general phenomenon in a rare-event and high-dimensional setting is the following: the true target distribution $\eta^{\ast}_{l}$ for $l$ large is extremely concentrated around very specific but unknown points in $\mathcal{X}$, so that, unless the pool of candidates among which snapshots $(X_{1},\ldots,X_{k})$ are picked is partly also located in such regions, the information given by the reduced model may eventually lack the accuracy required in order to describe the rare event. As a consequence, unless one is able to sample an extremely large amount of candidates, it is critical to select such candidates in areas that carefully track the true rare event. To make the picture clearer, consider at iteration $k$ the following surrogates of the target conditional distributions (1.1) built using the reduced score:
@@
where the normalization $Z_{l}\displaystyle\mathop{=}^{{\rm def}}\int_{\mathcal{X}}\,{\rm l}{\rm l}_{S>l}\mathrm{d}\pi,$ represents an approximation of the probability of the rare event. Then, for large $l$, the reduced rare event is also rare, hence very concentrated, and thus usually too far away from the true target rare event. If we want to guarantee sufficient mass in the region of the target rare event, we propose to specify a maximum possible level $l$ below which the true conditional distribution $\eta_{l}^{\ast}$ and the reduced one $\mu_{l}$ do match $\eta_{l}^{\ast}\simeq\mu_{l}$. This delicate issue is only rarely addressed in the literature. Most references avoid it and generate pools of candidates for the choice of snapshot relying on the reference distribution $\pi$. To the best of our knowledge, the only related pieces of work deal with the reverse problem of selecting (in a hierarchy of pre-computed surrogates) [16] or refining [3] the reduced score $S$, so that it is appropriate for a certain level $l$ according to some error quantification.
The main contribution of the present work is to use the following guidelines to design an efficient and adaptive algorithm that synchronizes: i) the sequential Monte Carlo sampling of rare events using a reduced model (that is sampling the flow $l\mapsto\mu_{l}$), and ii) the sequential update of the reduced model with evaluations of the true model.
Reduced Monte-Carlo simulation
Strategies for Monte-Carlo sampling of the reduced conditional distributions $\mu_{l}$ can be grouped in two broad categories: on the one hand importance sampling (IS), usually through a parametric approach, and on the other hand splitting methods, i.e., sequential Monte Carlo [11] methods, which are non-parametric, and in particular AMS [5] also called subset simulation [1].
Many approaches to IS of the reduced distribution $\mu_{l}$ have been proposed in recent years. In particular, cross-entropy methods [10] have been proposed relying on kriging [2] or reduced basis approximations [14, 16]. In order to increase the effectiveness of the cross-entropy method, it has also been suggested to approximate the parameter variable $X$ by constraining its support in well-chosen low dimensional subspaces of $\mathcal{X}$ [25]. Other recent contributions aim to build specific adaptive proposals for IS [23, 27] or for stratified sampling [4] (that can be interpreted as a form of IS). However, the effectiveness of IS approaches depends heavily on the choice of parametric family of proposals, for which we don’t necessarily have any prior knowledge.
On the other hand, splitting methods have also been used in conjunction with (adaptively constructed) kriging-based surrogate models [18, 3]. We refer again to the review on methods related to active learning and kriging provided in [20]. However, in these works, the splitting method samples an approximation $\mu_{l}$ of the target probability (1.1), which may be biased. We note nevertheless that in [13] subset simulation is performed using nested multiresolution sets (provided by an accuracy hierarchy of the surrogate model), yielding an estimate with an associated a posteriori error.
Adapting splitting methods to reduced scores can be optionally supplemented with a bridging procedure. The goal of this bridge is to sample a new proposal $\mu^{\prime}_{l^{\prime}}$ defined with some updated reduced score ${S}^{\prime}$ and level $\l^{\prime}$, starting from the samples of the current proposal $\mu_{l}$. In general, this sampling is not trivial. For example, the nesting underlying a AMS simulation is broken, since in general updating the reduced score and the level will imply $\left\{{S}^{\prime}>l^{\prime}\right\}\not\subset\left\{S>l\right\}$. There is very little work in the literature on this bridging problem. Let us mention the work [26], which proposes a bridging procedure with a constant $l^{\prime}=l$ level. Specifically, the authors use the sequential Monte Carlo method with soft levels and tempering instead of targeting distributions defined by conditioning on surrogate rare events. The authors mention that such an approach outperforms the multilevel subset simulation [24] algorithm they previously proposed, which dealt with the nesting problem by computing correction terms and performing Markov chain Monte Carlo simulations.
1.3 Contributions
In this work, we propose a sequential algorithm relying on reduced scores and error bounds. The algorithm provides an answer to the questions and limitations listed in Section 1.2. We consider the family of proposal distributions (1.3) (built using the reduced score) for sampling the target conditional distributions (1.1). The sketch of the proposed algorithm is the following. At each iteration $k$, the algorithm
  • computes adaptively a maximum level denoted $l^{(k)}$ such that the proposal $\mu_{l^{(k)}}^{(k)}$ matches the true conditional distribution $\eta^{\ast}_{l^{(k)}}$, as enabled by the accuracy of the current reduced model $S^{(k)}$
  • simulates (with a population Monte Carlo procedure) the proposal $\mu_{l}^{(k)}$ from an initial level $l=l_{b}^{(k)}$ to the maximum level $l=l^{(k)}$;
  • samples the new snapshot $(X^{(k)},\Psi^{\ast}(X^{(k)}))$ according to a weighted version of the distribution $\mu_{l^{(k)}}^{(k)}$. The weights (or learning function) may favor states $x$ with large reduced score error $E^{(k)}(x)$ (see (1.2)).
  • updates the reduced model score $S^{(k)}\to S^{(k+1)}$.
  • computes (using records of previous iterations) the next initial level ${l_{b}^{(k+1)}}$ so that the nesting (inclusion) of the support of the considered proposals are guaranteed.
  • bridges with (population) Monte Carlo simulation records of previous iterations to the next proposal $\mu_{l_{b}^{(k+1)}}^{(k+1)}$.